OpenStudy (anonymous):

Mrs. Flores gave the waiter $45 for her dinner that included a 25% tip. What is the cost of her dinner before tip?

OpenStudy (anonymous):

I know the answer is $36. I have no idea how to work the problem to get to that answer.

OpenStudy (anonymous):

I got it! 45 / 1.25=$36

OpenStudy (anonymous):

Cost of the meal is x so you pay for the meal which is x and you pay for the tip which is .25x so x + .25x = 45 1.25x = 45 x = 45/1.25 x = 36